OpenStudy (anonymous):

how do you find the x and y intercept of f(x)= (1)/(x+2)

OpenStudy (whpalmer4):

x intercept: f(x) = 0, find x that makes f(x) = 0. y intercept: f(0) = ?, find value of f(0)

OpenStudy (anonymous):

can someone explain further

OpenStudy (whpalmer4):

There's not much to explain. \[f(x) = \frac{1}{x+2}=0\] Is there a value of x for which that is true? No, not even x = a zillion kajillion :-) No x intercept. y intercept is another story. \[f(0) = \frac{1}{0+2} = \frac{1}{2}\] So the y intercept is 1/2.

OpenStudy (anonymous):

okay thankyou

OpenStudy (whpalmer4):

You can see that the y-intercept is 1/2, and that it doesn't look like it will ever cross the x-axis...